what is next in the sequence 
16,8,8,12
a)12  b)24  c)30  d)32

Grade:10

2 Answers

ekta kumar varma
8 years ago
12
Hasan Naqvi
8 years ago
16*0.5= 8
8 * 1  = 8
8 *1.5= 12
12 * 2 = 24
 
Ans = b)24